OpenStudy (anonymous):

Solve this exponential equation: 5(4^3y-5)=40

OpenStudy (anonymous):

if your solving for why the answer would be y=13/64 I believe

ganeshie8 (ganeshie8):

start by dividing 5 both sides

OpenStudy (anonymous):

The first step is to divide each side by 5 This will give you 4^3y-5 = 8

OpenStudy (anonymous):

You can then add 5 to each side This will give you 4^3y = 13

OpenStudy (anonymous):

The power of 3 is very tricky. You will want to cube root both sides to get rid of it. This will give you 4y = cuberoot(13) This may be the required answer because simplifying this is going to have a lot of decimals and get really messy. Does this make sense?

OpenStudy (anonymous):

The answer should be 13/6. I don't understand why I would cuberoot @KirbyLegs. Why can't I just take the log of both sides?

OpenStudy (anonymous):

she tooooooold you ;x

OpenStudy (anonymous):

LOL

hartnn (hartnn):

the question is actually

OpenStudy (anonymous):

When i did the solution, y=5.410459404 , which is clearly the wrong answer. I don't know what I did wrong :$

hartnn (hartnn):